How many gallons of a 30 percent alcohol solution must be mixed with 10 gallons of a 50 percent alcohol solution to make a 40 percent alcohol solution? The equation is ((( Amount x Concentration = Quantity)))
Answer by Paul(988)


Since 30% = 30x
Since 10 gallon of 50% = 50(10-x)
Total = 40(10)
Add it together:
30x+50(10-x)=40(10)
30x+500-50x=400
-20x=-100
x = 5
Hence, there should be about 5 gallons of 30%.
Paul.